OpenStudy (anonymous):

integral of sqrt(x)/(1+x^3)

OpenStudy (anonymous):

u=x^3/2

OpenStudy (anonymous):

i was gonna try using integration by parts

OpenStudy (anonymous):

but that isnt working

OpenStudy (anonymous):

did you use the u substitution?

OpenStudy (anonymous):

how did you know to use u=x^(3/2)?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

diff of x^(3/2) is 3/2 * sqrt(x)

OpenStudy (anonymous):

\[\int\frac{\sqrt x}{1+x^3}~dx\] Let \(u=x^{3/2}~\Rightarrow~du=\dfrac{3}{2}x^{1/2}~dx\). (i.e. \(\dfrac{2}{3}du=\sqrt x~dx\).) \[\frac{2}{3}\int\frac{du}{1+\left(u^{2/3}\right)^3}\\ \frac{2}{3}\int\frac{du}{1+u^2}\] Carry on with a trig sub.

OpenStudy (anonymous):

thanks. Its hard to figure out where to start sometimes with integrals, but im starting to figure out patterns as to which technique to use for certain forms of them

OpenStudy (anonymous):

you're welcome!

OpenStudy (jhannybean):

Now can you explain WHY u =x^(3/2)?

OpenStudy (anonymous):

because x^2 raised to the 3/2 power gives me x^3.

OpenStudy (jhannybean):

Oh I see... you're trying to rework it into the definition of tangent inverse... \(\large \int \cfrac{1}{1+x^2}dx\)

OpenStudy (anonymous):

yes, something like that. :)

OpenStudy (anonymous):

but i am ultimately trying to use trig substitution to get it into that form.

OpenStudy (jhannybean):

Are you still stuck?

OpenStudy (anonymous):

i figured it out actually. Thanks though :)

OpenStudy (jhannybean):

\begin{align} \large \int\frac{\sqrt x}{1+x^3}~dx &\\ \\ & \text{let u}= x^{3/2} \iff x=u^{2/3} \ , \ du = \frac{3}{2} x^{1/2}dx \rightarrow \frac{2}{3}du = \sqrt{x} dx \\ \\ &= \large \cfrac{2}{3}\int \frac{1}{1+(u^{2/3})^3}du \\ &= \large \cfrac{2}{3}\int \frac{1}{1+u^{(2/3) \cdot 3}} dx \\ &= \large \cfrac{2}{3}\int \frac{1}{1+u^2}du \\ &= \large \cfrac{2}{3} \tan^{-1}(u) +c \\ &= \large \frac{2}{3} \tan^{-1} (x^{3/2}) +c \end{align}